OpenStudy (anonymous):

Evaluate when a=2, b=-3, c=1/2, and d=10

hartnn (hartnn):

whats the expression ?

OpenStudy (anonymous):

\[\left( a ^{b}b \right)^{3}\div a ^{2}bc ^{4}\]

OpenStudy (anonymous):

sorry for not writing the expression faster

hartnn (hartnn):

ok, lets take one by one a^b = 2^(-3) = ?

OpenStudy (anonymous):

.125

OpenStudy (anonymous):

\[(a ^{b}b)^{3}=(2^{-3}(-3))^{3}=(\frac{ 1 }{ 2^{3} }(-3))^{3}=(\frac{ -3 }{ 8 })^{3}\]

OpenStudy (anonymous):

\[a ^{2}bc ^{4}=2^{2}(-3)(\frac{ 1 }{ 2 }^{4})=\frac{ 2^{2}(-3)1^{4} }{ 2 }\]

OpenStudy (anonymous):

but in decimal form that = -.052734375

OpenStudy (anonymous):

I got .0197753906

OpenStudy (anonymous):

so, \[(\frac{ -3 }{ 8 })^{3}:\frac{ 2^{2}(-3)1^{4} }{ 2 }=\frac{ (-3)^{3} }{ 8^{3} }\times \frac{ 2 }{ 2^{2(-3)1^{4}} }\]

OpenStudy (anonymous):

\[\frac{ (-3)^{3} }{ 8^{3} }\times \frac{ 2 }{ 2^{2}(-3)1^{4} }\]

OpenStudy (anonymous):

no need to change it into decimal

OpenStudy (anonymous):

let it in fraction

OpenStudy (anonymous):

thx ☺

OpenStudy (anonymous):

but i got 81/2048?? is that even correct

OpenStudy (anonymous):

\[\frac{ -27 }{ 512 }\times \frac{ 2 }{-12 }\]

OpenStudy (anonymous):

can you solve it?

OpenStudy (anonymous):

9/1024

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

serously thats correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

u're welcome